OpenStudy (diamond77108):

1. Kim's age is twice that of her sister. When you add Kim's age to her sister's age, you get 36. How old is each sister?

OpenStudy (farmboy813):

18 i think

OpenStudy (farmboy813):

are there any answer choices

OpenStudy (amorfide):

kims age is twice that of her sister if her sisters age is a number, let's say, x and kim is twice that, so 2x then you add them together to get 36 2x+x=36 3x=36 x=12 so kim is 2x=2(12)=24, and her sister is 36-24=12

OpenStudy (diamond77108):

thx